OpenStudy (anonymous):

What is the radius of a circle with the equation ? x^2+y^2+6x+2y+3=0 Round your answer to the nearest thousandth.

OpenStudy (angela210793):

(x^2+6x+9)-9+(y^2+2y+1)-1+3=0 (x+3)^2+(y+1)^2=7